# Juntao Zhao1,*    Wenhao Lu2,†    Sheng Wang1,†    Lingpeng Kong1    Chuan Wu1

### v2

@@ 7   Impact Statement
This paper introduces QSpec, a novel quantization paradigm that synergistically combines complementary quantization schemes for speculative decoding to enhance computational efficiency while preserving model fidelity. The impact of QSpec is twofold.
From an academic perspective, QSpec establishes a new paradigm that decouples speed acceleration from quality preservation—a longstanding trade-off in prior quantization research. By enabling independent optimization of these objectives, our work addresses critical challenges in quantization design—particularly under modern training regimes where improved model convergence and higher parameter utilization efficiency amplify the severity of precision degradation caused by quantization  Kumar et al. (2024).
@@
For industry applications, QSpec provides a practical pathway to accelerate inference without compromising output quality through efficient low-precision kernels. This advancement challenges hardware vendors to reconsider architectural support for low-precision execution, including specialized instruction set architectures (ISAs) and memory subsystems. The plug-and-play nature of QSpec further facilitates seamless integration into existing deployments of quantized models on edge devices, lowering barriers to adoption.
We identify no direct adverse societal impacts, as our work focuses on optimizing computational efficiency for language model serving systems. By reducing resource requirements, QSpec could broaden access to high-performance AI infrastructure, thereby contributing to more sustainable and scalable deployments of language technologies.
@@ A.1   Acceleration Evaluation
Versus quantization configurations.
Table 7: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. “Avg.” denotes the average speedup ratio for the corresponding row or column. $\dagger$” indicates the failure of W4A16 kernels to support these batch sizes together with long sequences and the large models.
@@
QSpec 8 594.1 (1.41$\times$) 648.2 (1.36$\times$) 760.1 (1.26$\times$) 723.6 (1.35$\times$) 787.5 (1.29$\times$) 738.8 (1.32$\times$) 1.33$\times$
16 811.5 (1.40$\times$) 936.0 (1.31$\times$) 1157.8 (1.17$\times$) 1042.1 (1.30$\times$) 1294.5 (1.20$\times$) 1171.4 (1.27$\times$) 1.27$\times$
32 1030.4 (1.42$\times$) 1240.2 (1.33$\times$) 1617.4 (1.05$\times$) 1248.5 (1.31$\times$) 1969.6 (1.16$\times$) 1576.0 (1.18$\times$) 1.24$\times$
Avg. 1.41$\times$ 1.33$\times$ 1.16$\times$ 1.32$\times$ 1.21$\times$ 1.25 $\times$ 1.28$\times$
@@
QSpec 8 253.7 (1.54$\times$) 291.5 (1.51$\times$) 298.3 (1.33$\times$) 350.9 (1.46$\times$) 345.7 (1.42$\times$) 310.3 (1.41$\times$) 1.44$\times$
16 359.8 (1.55$\times$) 420.2 (1.47$\times$) 466.7 (1.21$\times$) 555.2 (1.36$\times$) 557.8 (1.28$\times$) 473.1 (1.32$\times$) 1.37$\times$
32 441.8 (1.64$\times$) 527.2 (1.46$\times$) 575.3 (1.20$\times$) 749.4 (1.35$\times$) 770.0 (1.24$\times$) 628.4 (1.34$\times$) 1.39$\times$
Avg. 1.58$\times$ 1.48$\times$ 1.25$\times$ 1.39$\times$ 1.31$\times$ 1.36$\times$ 1.39$\times$
@@
32 $\dagger$ $\dagger$ 509.8 468.7 706.0 580.5
QSpec 8 222.6 (1.43$\times$) 233.9 (1.35$\times$) 256.7 (1.19$\times$) 271.5 (1.30$\times$) 285.0 (1.23$\times$) 268.3 (1.24$\times$) 1.29$\times$
16 322.6 (1.45$\times$) 362.5 (1.38$\times$) 402.7 (1.14$\times$) 438.5 (1.27$\times$) 507.5 (1.20$\times$) 453.5 (1.23$\times$) 1.28$\times$
32 400.2 ($\dagger$) 362.5 ($\dagger$) 578.1 (1.13$\times$) 573.0 (1.22$\times$) 798.8 (1.13$\times$) 684.5 (1.18$\times$) 1.27$\times$
Avg. 1.44$\times$ 1.36$\times$ 1.15$\times$ 1.26$\times$ 1.19$\times$ 1.22$\times$ 1.27 $\times$
@@
32 $\dagger$ $\dagger$ 376.4 365.5 255.9 287.4
QSpec 8 148.2 (1.56$\times$) 167.9 (1.49$\times$) 193.6 (1.35$\times$) 201.2 (1.44$\times$) 194.5 (1.33$\times$) 174.0 (1.36$\times$) 1.42$\times$
16 212.8 (1.56$\times$) 248.6 (1.45$\times$) 316.8 (1.26$\times$) 323.3 (1.36$\times$) 327.4 (1.28$\times$) 266.9 (1.29$\times$) 1.29$\times$
32 266.6 ($\dagger$) 320.0 ($\dagger$) 451.5 (1.20$\times$) 483.0 (1.32$\times$) 327.4 (1.28$\times$) 379.3 (1.32$\times$) 1.32$\times$
Avg. 1.56$\times$ 1.47$\times$ 1.27$\times$ 1.37$\times$ 1.29$\times$ 1.32$\times$ 1.38$\times$
Table 7 presents the comprehensive comparison of token generation throughput across multiple dimensions: model sizes, quantization configurations, and batch sizes, evaluated on various datasets. It is noteworthy that Llama2-7B shows higher speedup than Llama3-8B. This stems from the size difference primarily related to vocabulary, coupled with the introduction of Group-Query Attention (Ainslie et al., 2023), reducing the computation workload.
@@
Method Batch Size GSM8K (8-shot) MATH (4-shot) MBPP (0-shot) HumanEval (0-shot) ShareGPT LMsys-1k
EAGLE 1 65.81 70.11 68.53 49.15 79.60 71.29
@@ 16 OOM OOM OOM OOM OOM OOM
QSpec 8 208.95 (1.49$\times$) 249.97 (1.19$\times$) 185.13 (1.34$\times$) 185.99 (1.36$\times$) 329.44 (1.33$\times$) 260.48 (1.55$\times$)
16 292.82 356.93 269.87 255.11 562.07 463.35
@@ A.2   Ablation Studies
Comprehensive Performance Trade-offs: QSpec versus W4A16
We conduct a thorough analysis of the trade-offs between throughput and accuracy for our proposed framework against all baseline implementations. Figure 7 illustrates this comparison, plotting generation quality (accuracy) against computational efficiency (throughput). Our analysis reveals that while W4A4 suffers substantial performance degradation (18.5-39.5% reduction) on multi-step reasoning benchmarks compared to W4A16, QSpec achieves comparable accuracy to W4A16 while delivering significantly higher throughput. Although QSpec’s accuracy is marginally lower than W16A16 due to weight quantization-induced memory optimization, it successfully preserves the performance characteristics of W4A16 while offering superior computational efficiency.
@@ A.3   Datasets and Sampling
For the PIQA and Winogrande datasets, we randomly select 500 questions from each for performance evaluation. In contrast, we process the entire GSM8K and MATH datasets, as detailed in Atom Zhao et al. (2024b). When adapting QSpec to the Quarot method, we sample 200 samples from the GSM8K dataset, while sampling 700 from the MATH dataset, ensuring balanced representation by selecting 100 questions from each of the seven distinct question types. Additionally, we sample 200 questions from the MBPP dataset, while processing the entire HumanEval dataset. This preserves a thorough assessment of the model’s performance across various datasets.
For acceleration evaluation, we maintain the random seed at 42, and sample 100 samples from the GSM8K, MATH, MBPP, HumanEval, ShareGPT, and LMSys datasets, respsectively. This consistent methodology guarantees that our evaluation remains reproducible and representative across these diverse datasets.
@@ Appendix B Evaluation Datasets
  • LMsys-chat-1M-1K Zheng et al. (2023): Gathering one million authentic conversations with 25 leading large language models, this dataset was sourced from over 210,000 unique IP addresses interacting with the Vicuna demo and Chatbot Arena websites.

### v3

@@ 7.1 Impact Statement
This paper introduces QSpec, a novel quantization paradigm that synergies complementary quantization schemes through speculative decoding to enhance computational efficiency while preserving model fidelity. The impact of QSpec is twofold.
From an academic perspective, QSpec establishes a new paradigm that decouples efficiency from quality preservation—a longstanding trade-off in prior quantization research. This is achieved by the complementary quantization schemes: a low-precision activation-weight joint quantization for fast token drafting, and a high-precision weight-only quantization for accurate verification, enabling the independent optimization for efficiency and quality. This illuminates the pursuit of extreme efficiency in quantization schemes (e.g., W4A4) without the concern of performance degradation.
@@
For industry applications, QSpec provides a practical solution to accelerate inference without compromising output quality through efficient low-precision kernels. This prompts hardware vendors to reconsider their architectural support for low-precision execution, including specialized instruction set architectures (ISAs) and memory subsystems. Besides, the plug-and-play property of QSpec further facilitates seamless integration into existing deployments of quantized models in memory-constraint scenarios (e.g., edge devices).
Our research, focused on improving the computational efficiency of language model serving systems, is not anticipated to have direct negative social impact.
7.2 Limitation Discussion
The superior performance of QSpec relies on the high acceptance rate, particularly in small to moderate batch-size scenarios where the throughput gap between low- and high-precision quantization becomes pronounced. In contrast, traditional tree-based speculative decoding methods falter in batch serving, as discussed in Sec. 4, making QSpec’s advantages most evident in these settings. However, QSpec exhibits potential limitations in single-request scenarios, where other methods are preferentially optimized. With the popularity of LLMs and increasing batch serving, regular speculative decoding methods, including Medusa and EAGLE, degrade significantly (e.g., Medusa’s speedup drops below 1 at batch size of 16 in their Figure 22 Cai et al. (2024)), whereas QSpec excels.
To further address the limitations of QSpec in single-request scenarios, future research will focus on leveraging its high acceptance rate and reducing the overhead of the draft stage. Specifically, we aim to develop adaptive mechanisms that dynamically adjust the draft model’s sparsity, balancing latency and acceptance rate to achieve robust performance across both single-request and batch-size settings. Additionally, exploring hardware-aware optimizations, such as tailored low-precision kernels for resource-constrained devices, will enhance QSpec’s applicability in edge deployments. Furthermore, integrating QSpec into popular repositories (e.g., vLLM [^1] https://docs.vllm.ai/en/latest/) is also part of our future work. By unifying these advancements, QSpec can evolve into a versatile quantization framework, delivering consistent acceleration and fidelity for diverse inference scenarios, from personal devices to large-scale serving systems.
@@ A.1 Acceleration Evaluation
Table 6: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. “Avg.” denotes the average speedup ratio for the corresponding row or column.
@@
[…]
@@
[…]
@@
32 299.3 363.3 509.8 468.7 706.0 580.5
[…]
@@
32 207.5 241.6 376.4 365.5 255.9 287.4
[…]
Versus quantization configurations. Table 6 presents the comprehensive comparison of token generation throughput across multiple dimensions: model sizes, quantization configurations, and batch sizes, evaluated on various datasets. It is noteworthy that Llama2-7B shows higher speedup than Llama3-8B. This stems from the size difference primarily related to vocabulary, coupled with the introduction of Group-Query Attention (Ainslie et al., 2023), reducing the computation workload.
@@
Method
Batch Size GSM8K (8-shot) MATH (4-shot) MBPP (0-shot) HumanEval (0-shot) ShareGPT LMsys-1k
[…]
@@ 16 OOM OOM OOM OOM OOM OOM
[…]
@@ A.2 Ablation Studies
Performance Trade-offs: QSpec versus W4A16. We conduct a thorough analysis of the trade-offs between throughput and accuracy for our proposed framework against all baseline implementations. Figure 6 illustrates this comparison, plotting generation quality (accuracy) against computational efficiency (throughput). Our analysis reveals that while W4A4 suffers substantial performance degradation (18.5%-39.5% reduction) on multi-step reasoning benchmarks compared to W4A16, QSpec achieves comparable accuracy to W4A16 while delivering significantly higher throughput. Although QSpec’s accuracy is marginally lower than W16A16 due to weight quantization-induced memory optimization, it successfully preserves the performance characteristics of W4A16 while offering superior computational efficiency.
@@ A.3 Datasets and Sampling
For the PIQA and Winogrande datasets, we randomly select 500 questions from each for performance evaluation. In contrast, we process the entire GSM8K and MATH datasets, as detailed in Atom Zhao et al. (2024b). When adapting QSpec to the Quarot method, we sample 200 samples from the GSM8K dataset, while sampling 700 from the MATH dataset, ensuring balanced representation by selecting 100 questions from each of the seven distinct question types. Additionally, we sample 200 questions from the MBPP dataset, while processing the entire HumanEval dataset. This preserves a thorough assessment of the model’s performance across various datasets. For acceleration evaluation, we maintain the random seed at 42, and sample 100 samples from the GSM8K, MATH, MBPP, HumanEval, ShareGPT, and LMSys datasets, respsectively. This consistent methodology guarantees that our evaluation remains reproducible and representative across these diverse datasets.
A.4 QSpec on vLLM: Real-World Serving Evaluation
[…]
Our integration of QSpec into vLLM also initially aimed to enable a fair comparison with EAGLE (Li et al., 2024b). Unfortunately, we could not reproduce EAGLE’s performance on vLLM, as it exhibited significant performance degradation in batched scenarios, even losing speedup in the W4A16 setting. Moreover, due to the poor performance and inconsistent compatibility of other speculative decoding methods on vLLM, our additional experiments here do not include comparisons with other methods. We also note that the EAGLE team released EAGLE-3(Li et al., 2025) in late March 2025, open-sourcing partial pretrained model weights in April and claiming competitive speedup on vLLM. However, as of now, the code to reproduce this claim on vLLM is unavailable, preventing us from evaluating EAGLE-3’s performance on vLLM. Thus, we exclude comparisons with EAGLE-3, in line with the guidelines on handling contemporaneous work. This does not detract from the fact that both QSpec and EAGLE-3 represent outstanding contributions developed concurrently. As discussed in Section 7.2, QSpec and EAGLE exemplify two distinct design paradigms: EAGLE follows a training-intensive route with an independently optimized draft model and a refined acceptance policy, while QSpec adopts a training-free framework that leverages shared weights and KV cache in a unified structure. Each approach has its own strengths, and we look forward to future opportunities for dialogue and collaboration with the EAGLE team.
Table 8: Performance of QSpec on vLLM across different batch sizes and test sets, with acceptance rates per test set respectively. All speedup values are reported with a $\times$ multiplier.
Test Set Batch Size Acceptance Rate (%)
1 2 4 8 16 32
[…]
A.5 Artifact Documentation
We provide the official implementation of QSpec in the supplementary materials. The codebase is fully documented and includes [^4] https://github.com/hku-netexplo-lab/QSpec:
  • Installation guide covering dependency setup (CUDA 12.5, Python 3.10), environment recommendations (NVIDIA A100 or L20), and instructions for installing required third-party libraries and compiling QSpec kernels.
  • Docker support for reproducible deployment, including editable mount path and data path configurations, and build/run scripts.
  • Execution scripts for reproducing our throughput and latency results using demo.py, which supports different model paths, speculative token lengths, and batch sizes.
  • Pretrained models hosted on Huggingface for QSpec’s Llama-3-8b-instruct model used in our vLLM experiments.
  • Notes and caveats describing limitations of our current implementation (e.g., not optimized for all GPU types, partial vLLM integration, cold start auto-tuning delay).
  • License and Intended Use: We confirm that all third-party artifacts used in this work (e.g., vLLM, Huggingface-hosted models) were accessed and used in accordance with their licenses (Apache License 2.0) and intended research purposes. Our implementation of QSpec is released under the Apache License 2.0 and is explicitly intended for academic and non-commercial use. Users are instructed to obtain such resources directly from their original providers and to comply with the corresponding terms of use.
All artifacts are accompanied by a README.md file that details the usage and experimental instructions. The code is released under an anonymous GitHub repository to ensure reproducibility.
A.6 Understanding FP16 vs. W4A16 Performance in Main Results
While W4A16 quantization (e.g., AWQ) is often expected to outperform FP16 in small-to-medium batch sizes due to its design for improved efficiency in weight-only quantization Lin et al. (2024a), our main results consistently show FP16 surpassing W4A16 across various implementations, including vLLM, Atom’s system, and even Hugging Face’s official benchmarks555huggingface.co/docs/transformers/main/quantization/awq. This discrepancy may lead some readers to question the relative performance, as implementation-specific factors such as device characteristics, kernel optimizations, and system engineering significantly influence outcomes. We conducted a complementary experiment to elucidate this phenomenon, illustrated in Figure 7, comparing FP16 and W4A16 under different implementation settings.
In this experiment, Atom-FP16 and Atom-AWQ are derived from Atom’s system implementation, following the end-to-end benchmark settings of our main experiments (Section 4), with FlashInfer integrated into a Punica-style serving system to support continuous batching Yu et al. (2022). Conversely, Benchmark-FP16 and Benchmark-AWQ are sourced from the AutoAWQ repository, leveraging an optimized AWQ kernel and Flash-Attention, but employing a classic dummy benchmark method (directly invoking the model for the context length duration)666github.com/casper-hansen/AutoAWQ. Additionally, we include vLLM-FP16 and vLLM-AWQ, implemented in the vLLM framework (Commit 9a7c3a0, Appendix A.4), to provide a broader perspective.
Figure 7 reveals distinct performance trends across batch sizes of 8, 16, and 32. In Atom’s implementation, FP16 consistently outperforms AWQ across all batch sizes, aligning with our main results. However, in the AutoAWQ dummy benchmark, AWQ exhibits superior throughput, reversing the trend. In vLLM, AWQ slightly outperforms FP16 at a batch size of 8, but FP16 surpasses AWQ at batch sizes of 16 and 32. These variations underscore the impact of system implementation and kernel optimization on relative performance. Since our main experiments strictly adhere to Atom’s setup for fair and rigorous comparisons (Sec. 4), the observed speed disparity between FP16 and W4A16 due to implementation differences does not undermine the validity of our claims regarding QSpec’s performance.
Figure 7: Normalized throughput of FP16 and AWQ implementations across batch sizes 8, 16, and 32. Atom-FP16 and Atom-AWQ use Atom’s system with FlashInfer and continuous batching. Benchmark-FP16 and Benchmark-AWQ are from AutoAWQ with optimized AWQ kernels and Flash-Attention, using a dummy benchmark method. vLLM-FP16 and vLLM-AWQ are implemented in vLLM. The generation length is set to 512.
@@ Appendix B Evaluation Datasets
  • LMsys-chat-1M-1K Zheng et al. (2023a): Gathering one million authentic conversations with 25 leading large language models (LLMs), this dataset was sourced from over 210,000 unique IP addresses interacting with the Vicuna demo and Chatbot Arena websites.
@@ Appendix C Datasets Sampling
Appendix D Related Work
Speculative Decoding leverages a draft model to generate candidate tokens, which are then validated by a target model (Leviathan et al., 2023). Recent research has primarily focused on improving the acceptance rate and generation speed of candidate tokens. SpecInfer (Miao et al., 2024) introduces a boost-tuned small language model to generate candidate tokens in tree structures, enabling single-pass verification. In contrast, EAGLE (Li et al., 2024b) adopts an aggressive pruning strategy for the draft model’s architecture, allowing penultimate layer feature prediction with minimal computational overhead. Self-speculative decoding, a subset of this technique, employs a single model for both draft generation and verification. LayerSkip (Elhoushi et al., 2024) introduces a training methodology for early exit with layer drop, subsequently verifying partially generated tokens through full model inference. Medusa (Cai et al., 2024) augments the original LLM with additional heads atop the final hidden state while relaxing the acceptance policy. However, these approaches inevitably require retraining of the original model, which can be computationally expensive and time-consuming. We further demonstrate their deficiency in batched serving under quantization scenario.
Tree-Structured Drafting. Tree-structured drafting (Chen et al., 2024; Miao et al., 2024; Li et al., 2024b, a) is a widely adopted technique to improve acceptance rates in speculative decoding. Instead of generating a single draft token at a time using the draft model $\mathcal{M}_{d}$, $\mathcal{M}_{d}$ select $k$ (i.e., top-k) tokens and infers $\gamma$ times to form a draft tree with depth $\gamma$. The target model $\mathcal{M}$ then verifies the tokens using masked tree attention, with the exact verification strategy depending on the sampling method (Cai et al., 2024; Miao et al., 2024). Under greedy sampling, the highest-probability token at each position is selected, forming the longest branch with the common prefix as the accepted sequence. In this tree, the root node presents the past sequence, any token $t_{i}$ has a path from the root node $r$ to $t_{i}$, denoted by $Path(r,t_{i})$, consisting of ancestors $a_{1},\ldots,a_{j}$.
Appendix E Supplementary Figures and Tables
To improve the clarity of the main text and streamline presentation, we provide additional visualizations and ablation results related to QSpec in this appendix. These include supporting data referenced in the main body and additional experiments.
Quantization Method
ShareGPT MATH (4-shot) MBPP (0-shot)
Atom 83.8 89.4 88.6
QuaRot 81.6 88.9 85.4
Table 10: Ablation study comparing acceptance rates (%) across large reasoning model and difficult reasoning tasks using QSpec.
Model GPQA-Diamond Super-GPQA AIME ARC MMLU
Llama3-8b
-Instruct
94.1 96.5 96.1 92.6 92.4
OpenBookQA RACE SQuAD v2 TruthfulQA HellaSwag
92.6 94.2 95.0 92.0 91.7
HumanEval LAMBADA Social IQa Avg.
87.5 89.6 93.8 92.93
DeepSeek-R1-Distilled
-QWen14B
96.2 96.0 97.9 90.4 90.8
OpenBookQA RACE SQuAD v2 TruthfulQA HellaSwag
90.5 92.8 92.8 88.6 96.7
HumanEval LAMBADA Social IQa Avg.
96.7 94.3 91.7 93.49
Table 11: Performance results for the reasoning model DeepSeek-R1-Distilled-QWen14B. We follow the settings of our main experiments and set the batch size (bs) to 16.
Dataset GSM8K MATH MBPP HUMANEVAL SHAREGPT LMSYS AVG.
W4A16 (tokens/s) 139.38 203.52 200.84 216.54 194.31 194.22 191.47
QSpec (tokens/s) 171.56 282.17 278.74 292.75 259.48 246.07 255.13
Speedup 1.23 1.39 1.39 1.35 1.34 1.27 1.33
Appendix F AI Assistance Statement
We used AI tools (e.g., ChatGPT) exclusively for language polishing and LaTeX formatting. No part of the core research, including ideation, experimental design, or analysis, was generated by AI tools. All scientific contributions are the sole work of the authors.